OpenStudy (anonymous):

\[y''+2y'+y=xe^{-x}\] \[r^2+2r+1=0\] \[y_c = c_1 e^{-x} + c_2 x e^{-x}\] \[y_p=Ae^{-x}+Bxe^{-x}\] \[y_p'=-Ae^{-x}-Be^{-x}(x-1)\] \[y_p''=Ae^{-x}+Be^{-x}(x-2)\] something is wrong here because I get x=0

OpenStudy (anonymous):

someone said I should have a Cx there somewhere....something something...didn't quite get it

OpenStudy (anonymous):

"There's a quantitative way to justify it, but I can't recall how it goes. The point is xe^-x is the forcing term and one of the solutions to the homogeneous. So you need another term in your yp if you're ever going to get an actual solution that satisfies the original non-homogeneous equation."

OpenStudy (turingtest):

yes you are missing an extra x it seems

OpenStudy (turingtest):

normally the particular for the non-homogeneous problem you have here would be\[Y_p=(Ax+B)e^{-x}\]but you must remember that all your solutions must remain linearly independent from each other. since your guess for the particular already appears in the complimentary, you must multiply by another x, making your guess for the particular\[Y_p=x(Ax+B)e^{-x}\]that should fix your problem

OpenStudy (anonymous):

so there is no C term?

OpenStudy (turingtest):

the problem is the \(c_2xe^{-x}\) term in the complimentary is the same as the \(xe^{-x}\) in the particular and no, there is no C term

OpenStudy (anonymous):

ok I'll try it again, let's see...

OpenStudy (anonymous):

\[y_p=Axe^{-x}+Bxe^{-x}\] \[y_p'=-Ae^{-x}(x-1)-Be^{-x}(x-1)\] \[y_p''=Ae^{-x}(x-2)+Be^{-x}(x-2)\] \[(Ae^{-x}(x-2)+Be^{-x}(x-2))+2(-Ae^{-x}(x-1)-Be^{-x}(x-1))+\]\[(Axe^{-x}+Bxe^{-x})=xe^{-x}\]

OpenStudy (anonymous):

I somehow cancel out all of the x-terms...I guess I'm doing something wrong

OpenStudy (anonymous):

oh I see what I did, let's try again

OpenStudy (anonymous):

I did the distributive property wrong

OpenStudy (turingtest):

...and you missed an x in the guess...

OpenStudy (turingtest):

\[y_p=(Ax^2+Bx)e^{-x}\]or is that what you meant?

OpenStudy (anonymous):

doesn't it disappear in the first and second derivative?

OpenStudy (turingtest):

but then where is the 2A ?

OpenStudy (anonymous):

oh there is an x^2, yeah that's what I did wrong

OpenStudy (turingtest):

yeah, one more time ;) good luck, I don't wanna do it on paper if I can help it :P

OpenStudy (anonymous):

yeah I should be fine =P thanks!